When I have a regression equation such as Y = beta 0 + beta 1 X and the estimated coefficients are beta hat 0 = 300, beta hat 1 = -4, SER = 10, and R^2 = .10, what is the regression's prediction of Y for an X of 20? What is Y bar for an X bar of 20.4?
Answer #1

The regression equation is y = 300 - 4*x

what is the regression's prediction of Y for an X of 20?

x = 20

y = 300 - 4*20 = 220

The prediction of Y for an X of 20 is 220.

b0 = 300

b1 = - 4